# Aerial Grasping via Maximizing Delta-Arm Workspace Utilization

### v1

@@ V-B1 Inclined object grasping
This scenario features fixed starting and ending points, with a stationary target point for the end-effector, positioned at $[0,0,1]$ m. To evaluate the aerial manipulator’s grasping performance under various attitude constraints, we established different inclination angles relative to the ground plane for the target point: $0^{\circ},10^{\circ},20^{\circ},30^{\circ},40^{\circ},50^{\circ},60^{\circ}$. When passing through the target point, the end-effector must reach the point and satisfy these angular requirements. We record the execution time and end-effector range of motion for both methods across different angles. The results are presented in Fig. 5.
As shown in Fig. 5 (a), our method demonstrates significant advantages in execution efficiency, with substantially lower completion times compared to [8]. Furthermore, due to the utilization of an expanded workspace, our approach enables a greater range of motion for the delta arm, with notably larger Z-axis activity scope in $\mathcal{F}_{A}$ than the baseline (c.f. Fig. 5 (b)). This enhanced operational range contributes to the observed performance improvements.
@@ V-C Real-World Experiments
We further validated the feasibility of our method in real-world experiments, as shown in Fig. 7. Our real-world platform consists of a quadrotor using NVIDIA Jetson Orin NX 16GB as the onboard computer and a delta arm with a hook-and-loop mechanism attached to its end-effector. The platform’s state estimation is achieved by fusing data from the NOKOV Motion Capture System and the inertia measurement unit (IMU). In this experiment, the aerial manipulator needs to execute an aerial grasping task, using the hook-and-loop mechanism to pick up rectangular boxes with corresponding hook-and-loop surfaces. The box is placed on a table of a certain height, and the experiment is divided into two scenarios (Experiment 1 and Experiment 2) based on different box placement angles. Due to the limited sensing capabilities of our experimental platform, we represented the scene using pre-constructed point cloud maps. The starting and ending points of the aerial manipulator remain fixed, while the position and orientation of the boxes are obtained through NOKOV Motion Capture System.
@@
As illustrated in Fig. 7, the aerial manipulator successfully completed the grasping tasks in both experiments, adhering to the target objects and transporting them away with speeds up to $3$ m/s. In Experiment 1, it performs high-speed grasping, picking up the object located $4$ meters away in just $1.8$ seconds. Fig. 8 (a-b) depicts the end-effector velocity and its position relative to the target grasping point throughout the task execution of the two methods, respectively. When approaching the target point, the end-effector velocity rapidly decreases under the influence of velocity constraints in the task part, reaching $0.49$ m/s and $0.40$ m/s at the grasping points respectively, ensuring stability during capture. Additionally, we assign relatively low weights to the velocity constraints to maintain reasonable operational efficiency. In terms of positioning accuracy, the end-effector achieves distances of $0.02$ m and $0.04$ m from the designated grasping points. These minimal errors ensured stable grasping while demonstrating the effectiveness of our framework.
@@ VI Conclusion
In this work, we present a whole-body grasp planning framework for aerial manipulators equipped with delta arms. We use a learning-based method to represent the complex delta arm workspace, maximizing its utility. Then, we formulate a coupled optimization problem to generate the trajectory for both the quadrotor and the delta arm, incorporating task constraints to ensure manipulation stability and effectiveness. We develop a learning-based approximation of the delta arm’s forward kinematics, enabling efficient workspace constraint elimination. The framework’s efficacy is validated through extensive simulations and real-world experiments. We show that our framework enables the UAV to grasp an object located 4 meters away within 2 seconds. Future improvements for our framework might include enhancing real-time performance and autonomous exploration capabilities.

### v2

@@ V-B1 Inclined object grasping
This scenario features fixed starting and ending points, with a stationary target point for the end-effector, positioned at [0,0,1][0,0,1][ 0 , 0 , 1 ] m. To evaluate the aerial manipulator’s grasping performance under various attitude constraints, we established different inclination angles relative to the ground plane for the target point: $0^{\circ},10^{\circ},20^{\circ},30^{\circ},40^{\circ},50^{\circ},60^{\circ}$. When passing through the target point, the end-effector must reach the point and satisfy these angular requirements. We record the execution time and end-effector range of motion for both methods across different angles. The results are presented in Fig. 5.
As shown in Fig. 5 (a), our method demonstrates significant advantages in execution efficiency, with substantially lower completion times compared to [8]. Furthermore, due to the utilization of an expanded workspace, our approach enables a greater range of motion for the delta arm, with notably larger Z-axis activity scope in $\mathcal{F}_{A}$ than the baseline (cf. Fig. 5 (b)). This enhanced operational range contributes to the observed performance improvements.
@@ V-C Real-World Experiments
We further validated the feasibility of our method in real-world experiments, as shown in Fig. 7. Our real-world platform consists of a quadrotor using NVIDIA Jetson Orin NX 16GB as the onboard computer and a delta arm with a hook-and-loop mechanism attached to its end-effector. The platform’s state estimation is achieved by fusing data from the NOKOV Motion Capture System and the inertia measurement unit (IMU). In this experiment, the aerial manipulator needs to execute an aerial grasping task, using the hook-and-loop mechanism to pick up rectangular boxes with corresponding hook-and-loop surfaces. The box is placed on a table of a certain height, and the experiment is divided into two scenarios (Experiment 1 and Experiment 2) based on different box placement angles. Due to the limited sensing capabilities of our experimental platform, we represented the scene using pre-constructed point cloud maps. The starting and ending points of the aerial manipulator remain fixed, while the position and orientation of the boxes are obtained through the NOKOV Motion Capture System. The controller design for the aerial manipulator follows the approach outlined in [20].
@@
As illustrated in Fig. 7, the aerial manipulator successfully completed the grasping tasks in both experiments, adhering to the target objects and transporting them away with speeds up to 333 m/s. In Experiment 1, it performs high-speed grasping, picking up the object located 444 meters away in just 1.81.81.8 seconds. Fig. 8 (a-b) depicts the end-effector velocity and its position relative to the target grasping point throughout the task execution of the two methods, respectively. When approaching the target point, the end-effector velocity rapidly decreases under the influence of velocity constraints in the task part, reaching 0.490.490.49 m/s and 0.400.400.40 m/s at the grasping points respectively, ensuring stability during capture. Additionally, we assign relatively low weights to the velocity constraints to maintain reasonable operational efficiency. In terms of positioning accuracy, the end-effector achieves distances of 0.020.020.02 m and 0.040.040.04 m from the designated grasping points. These minimal errors ensured stable grasping while demonstrating the effectiveness of our framework.
@@ VI Conclusion
In this work, we present a whole-body grasp planning framework for aerial manipulators equipped with delta arms. We use a learning-based method to represent the complex delta arm workspace, maximizing its utility. Then, we formulate a coupled optimization problem to generate the trajectory for both the quadrotor and the delta arm, incorporating task constraints to ensure manipulation stability and effectiveness. We develop a learning-based approximation of the delta arm’s forward kinematics, enabling efficient workspace constraint elimination. The framework’s efficacy is validated through extensive simulations and real-world experiments. We show that our framework enables the UAV to grasp an object located 4 meters away within 2 seconds. Future improvements for our framework might include enhancing real-time performance and autonomous exploration capabilities. Although the proposed method successfully finds locally optimal feasible solutions, global optimality remains unguaranteed. Future research will explore methods to ensure global optimality.